OpenStudy (anonymous):

need help starting out: use the method of variation of parameters to determine the solution to the differential equation: y''' -3y'' + 4y = e^(2x)

OpenStudy (turingtest):

first thing is to find the complimentary solution

OpenStudy (anonymous):

so far i have....

OpenStudy (lgbasallote):

you can just use undetermined coefficients...right?

OpenStudy (turingtest):

it says variation of parameters, so...

OpenStudy (lgbasallote):

oh wait...it said to use variation specifically

OpenStudy (anonymous):

(D^3 -3D^2 +4)y = e^2x. that factored the (d+1)(d-2)^2 = e^2x

OpenStudy (lgbasallote):

ohhh so this is what you were asking...

OpenStudy (lgbasallote):

you have to change D into m first then factor like algebra

OpenStudy (anonymous):

yea

OpenStudy (lgbasallote):

and also make the right side = 0

OpenStudy (turingtest):

you first want the homogeneous solution, so all that business should be =0

OpenStudy (turingtest):

beat me too it

OpenStudy (lgbasallote):

\m/

OpenStudy (anonymous):

yes, then what do I do. what would (d-2)^2 become? e^(4x^2)?

OpenStudy (lgbasallote):

what are you doing? im getting confused..

OpenStudy (anonymous):

i am trying to find the fundamental solutions right now

OpenStudy (anonymous):

set*

OpenStudy (lgbasallote):

ohh you're solving for roots huh

OpenStudy (lgbasallote):

(d-2)^2 means root of 2 with multiplicity 2 right?

OpenStudy (anonymous):

or at least the general ssolution

OpenStudy (lgbasallote):

when you have a root that's like \[m = a^2\] the solution would be \[\large C_1 e^{ax} + C_2x e^{ax}\] does that help?

OpenStudy (anonymous):

so i would have 3 fundamental solutions: { e^-x , e^2x , and xe^2x}??

OpenStudy (lgbasallote):

dont forget the constants

OpenStudy (lgbasallote):

\[\huge y_c = c_1 e^{-x} + c_2 e^{2x} + c_3 x e^{2x}\]

OpenStudy (anonymous):

oh, well then I would have ......YUP! that's what i was about to type... but maybe not as pretty as that

OpenStudy (lgbasallote):

haha :p now do yp

OpenStudy (anonymous):

can i not use Kramer's rule now?

OpenStudy (anonymous):

derive twice then use matrices?

OpenStudy (anonymous):

[g(x)W[y2,y3](x)] / [W[y1,y2,y3](x)]

OpenStudy (lgbasallote):

they do that in variation of parameters? o.O all i know is you copy the yc and derive thrice (since third order differential equation) and then use matrices (but i just use systems of equations)

OpenStudy (anonymous):

let me try to derive 3x's then.... oye...

OpenStudy (lgbasallote):

but seriously...idk this "kramer's rule" you're saying..

OpenStudy (anonymous):

it's okay

OpenStudy (turingtest):

I guess you would use cramer's rule here, though I'm really only used to doing this for second-order problems

OpenStudy (lgbasallote):

that's why i was saying systems would be nicer :C

OpenStudy (anonymous):

.....i hate this pellet. it's waaay too confusing

OpenStudy (anonymous):

HA! pellet was automatically filled in for sh*t

OpenStudy (lgbasallote):

it takes time..especially if you're oriented with one method only

OpenStudy (anonymous):

well, my teacher tries to give us at least one or two methods. but he ends up rushing himself and getting the problem wrong himslef...